# A greedy heuristic for graph burning

### v1

@@ 4.1 A greedy $(1/2)$-approximation algorithm for CMCP
The classical greedy approximation algorithm for MCP iteratively selects the subset that covers as many uncovered elements as possible. This algorithm achieves the best possible ratio of $(1-\frac{1}{e})$ [13, 36, 14]. The greedy approximation algorithm for CMCP follows the same idea. Given $k$ clusters of subsets of a universe of elements, iteratively selects the subset that covers as many uncovered elements as possible among the $k$ clusters (See Algorithm 2). Selecting a subset from a cluster can make it impossible to cover some elements in the following iterations; we refer to these as discarded elements. Figure 2 shows an instance where $OPT/2$ elements are covered at the first iteration and $OPT/2$ are discarded. In fact, $1/2$ is the approximation ratio of this algorithm (See Lemma 1 and Theorem 3), and Figure 2 serves as a tight example.
@@
1:$k$ clusters of subsets, $C_{1},C_{2},...,C_{k}$.
2:A set containing one subset from each cluster.
3:$s,B\leftarrow\emptyset$
4:$I\leftarrow\{1,2,...,k\}$
5:while $I\not=\emptyset$ do
6:     $i\leftarrow\operatorname*{argmax}_{j\in I}\max_{S\in C_{j}}|S-B|$
7:     $I\leftarrow I-\{i\}$
8:     $A\leftarrow\operatorname*{argmax}_{S\in C_{i}}|S-B|$
@@
Figure 2: In this instance for CMCP, $k=2$, $C_{1}=\{S_{1},S_{2},S_{3}\}$, $C_{2}=\{S_{4}\}$. The optimal solution $\{S_{2},S_{4}\}$ covers $OPT=10$ squares. If the greedy approximation algorithm selects $S_{3}$ at the first iteration, the returned solution will be $\{S_{3},S_{4}\}$, which covers $OPT/2=5$ squares.
@@
At each iteration, the number of newly discarded elements cannot exceed the number of newly covered elements.
@@
A set $s_{i}$ with maximum coverage $|s_{i}-B|$ is selected at each iteration, where $B$ is the set of covered elements. Thus, the discarded elements, if any, must belong to a subset with coverage smaller than $|s_{i}-B|$. ∎
@@
The greedy approximation algorithm for CMCP has an approximation ratio of $1/2$.
@@
Let $OPT$ be the size of the optimal solution, let $x_{i}$ be the number of newly covered elements at iteration $i$, let $y_{i}$ be the number of newly discarded elements at iteration $i$, and let $z_{i}$ be the number of available elements (not yet covered nor discarded) at the end of iteration $i$. Since a subset with maximum coverage is selected at each iteration, it must cover at least $1/(k-(i-1))$ of the $z_{i-1}$ available elements; otherwise, $OPT$ would not be optimal (See Eq. (12) and Eq. (13)). By Lemma 1, no more than $x_{i}$ elements can be discarded (See Eq. (14)). Notice that $z_{0}=OPT$ and $z_{k}=0$ because all the available vertices are covered at the last iteration.
$x_{i}\geq\frac{z_{i-1}}{k-(i-1)}$ (12)
$z_{i}=z_{i-1}-x_{i}-y_{i}$ (13)
$y_{i}\leq x_{i}$ (14)
The proof is by contradiction. Let us assume that the algorithm returns a solution that covers less than $OPT/2$ elements, i.e.,
$\sum_{i=1}^{k}x_{i}<\frac{OPT}{2}\ \ .$ (15)
Combining Eq. (14) with Eq. (15), we get
$\sum_{i=1}^{k}y_{i}\leq\sum_{i=1}^{k}x_{i}<\frac{OPT}{2}\ \ .$ (16)
Then, combining Eq. (13) with Eq. (16) we get
$\sum_{i=1}^{k}z_{i}>\sum_{i=1}^{k}z_{i-1}-OPT\ \ .$ (17)
Notice that $\sum_{i=1}^{k}z_{i-1}-\sum_{i=1}^{k}z_{i}=z_{0}-z_{k}$. Since $z_{0}=OPT$ and $z_{k}=0$, Eq. (17) reduces to
$z_{0}-z_{k}=OPT<OPT\ \ ,$ (18)
which is a contradiction. Therefore, $\sum_{i=1}^{k}x_{i}\geq OPT/2$. Namely, the greedy approximation algorithm for CMCP has an approximation ratio of $1/2$. ∎
@@ 4.2 A greedy heuristic for GBP
Algorithm 3 shows the proposed deterministic greedy heuristic (Gr) for GBP, which implements the greedy approximation algorithm for CMCP. In a nutshell, starting from the clusters with larger subsets, i.e., from $k-1$ to 0 (line 3), it iteratively selects the subset that covers as many uncovered elements as possible (line 7). In more detail, Gr receives as input a graph $G=(V,E)$, a guess $k$ on $b(G)$, and the distance matrix $d$, which can be computed independently using breadth-first search. Then, Gr computes the closed neighborhood $N_{k-1}[v]$ of each $v\in V$. In other words, it computes the adjacency list of $G^{k-1}$ (line 2). Afterward, it iteratively selects $k$ subsets that cover as many uncovered vertices as possible in graphs $G^{r}$, where $r$ goes from $k-1$ to 00 (lines 3 to 10). Notice that the adjacency list is updated at each iteration, containing uncovered vertices at a distance at most $r$ from each $v\in V$ (line 5). Besides, every subset $N_{i}[v]$ of each cluster is a superset of each $N_{j}[v]$, where $j<i$. Thus, a subset with maximum coverage among all the clusters is selected at each iteration. Therefore, this is a correct implementation of Algorithm 2. Since the distance matrix $d$ is known, updating the adjacency list can be done in at most $\mathcal{O}(n^{2})$ steps at each iteration. Thus, the overall complexity of Gr is $\mathcal{O}(kn^{2})$; however, if we consider the time required to compute all-pairs’ shortest path, its complexity is $\mathcal{O}(mn+kn^{2})$ as stated in the abstract. By Theorem 3, Proposition 4 follows.
@@
1:A graph $G=(V,E)$, its distance matrix $d$, and a guess $k$ on $b(G)$
2:A sequence
3:$s,B\leftarrow\emptyset$
4:$N_{k-1}\leftarrow$ Adjacency list of $G^{k-1}$
5:for $r=k-1$ to 00 do
6:     if $r<k-1$ then
7:         $N_{r}\leftarrow$ Update the adjacency list for $G^{r}$
8:     end if
9:     $v\leftarrow\operatorname*{argmax}_{u\in V}|N_{r}[u]-B|$
10:     $s\leftarrow s\cup\{(k-r,v)\}$
11:     $B\leftarrow B\cup N_{r}[v]$
12:end for
13:return $s$
@@
If the subsets Gr selects cover less than $n/2$ vertices, then $k<b(G)$.
@@
By Theorem 3, Algorithm 2 returns a solution that covers at least $OPT/2$ elements, where $OPT$ is the optimal solution size. By definition, a burning sequence burns $n$ vertices. So, if $k\geq b(G)$, $OPT=n$, and since Gr correctly implements Algorithm 2, it must return a solution that covers at least $n/2$ vertices. By contrapositive, if Gr returns a solution that covers less than $n/2$ vertices, then $k<b(G)$. ∎
For Gr to be useful, we must know $b(G)$. We can try different guesses $k$ to tackle the issue of not knowing this value in advance. The simpler way is by trying all $k\in[1,n]$. However, a binary search is more efficient. Although Proposition 4 helps find a lower bound for $b(G)$, we didn’t find it useful in practice. Therefore we didn’t integrate it into the binary search (See Algorithm 4), which we now explain. First, the distance between every pair of vertices is computed in $\mathcal{O}(mn+n^{2})$ steps using breadth-first search (line 1). Then, the lower and upper bounds are computed using the burning farthest-first (BFF) approximation algorithm for GBP [12], which runs in $\mathcal{O}(n^{2})$ steps if the computation of all-pairs’ shortest path is not considered (Algorithm 5 shows a detailed pseudocode for BFF using a FIFO queue). Since BFF returns a solution $s_{0}$ (line 2) of length at most $23b(G)-$ 2 [12], the lower bound on $b(G)$ can be safely set to $\lceil(len(s_{0})+2)/3\rceil$ (line 3). Then, Gr is executed with parameters $G$, $k$, and $d$ (lines 6 and 7); GrP will be introduced later, and if an exact solver is used, this is equivalent to Algorithm 1 using tighter bounds. Afterward, every time Gr returns a burning sequence, $h$ is set to $k-1$; otherwise, $l$ is set to $k+1$. This way, Gr is executed up to $\log n$ times with different guesses on $k$ that might approach $b(G)$. Using Gr at line 7, and since $k\leq n$, the complexity of Algorithm 4 is $\mathcal{O}(n^{3}\log n)$.
@@
5:$l\leftarrow\lceil(len(s_{0})+2)/3\rceil$
6:$h\leftarrow len(s_{0})$
@@
8:     $k\leftarrow\lfloor(l+h)/2\rfloor$
9:     $s\leftarrow$ Gr$(G,k,d)$ or GrP$(G,k,d)$ or ExactSolver$(\text{CMCP}(G,k,d))$
@@
2:A burning sequence of length at most $23b(G)-$ 2
@@
5:$i\leftarrow 1$
6:$s\leftarrow\{(i,v)\}$
7:$q.push(v)$
8:$B\leftarrow\{v\}$
9:for $u\in V$ do
10:     $dist(u)\leftarrow d(u,v)$
11:end for
12:while $B\not=V$ do
13:     $q\_size\leftarrow q.size()$
14:     for $j=1$ to $q\_size$ do
15:         $v\leftarrow q.pop()$
16:         for $u\in N(v)-B$ do
17:              $q.push(u)$
18:              $B\leftarrow B\cup\{u\}$
@@
20:     end for
21:     $v\leftarrow\operatorname*{argmax}_{u\in V}dist(u)$
22:     $i\leftarrow i+1$
23:     $s\leftarrow s\cup\{(i,v)\}$
24:     $q.push(v)$
25:     $B\leftarrow B\cup\{v\}$
26:     for $u\in V$ do
27:         if $d(u,v)<dist(u)$ then
28:              $dist(u)\leftarrow d(u,v)$
29:         end if
30:     end for
31:end while
32:return $s$

### v3

@@ 4.1 A greedy $(1/2)$-approximation algorithm for CMCP
The classical greedy approximation algorithm for MCP iteratively selects the subset that covers as many uncovered elements as possible. This algorithm achieves the best possible ratio of $(1-\frac{1}{e})$ [13, 18, 14]. The greedy approximation algorithm for CMCP follows the same idea. Given $p$ clusters of subsets of a universe of elements, iteratively selects the subset that covers as many uncovered elements as possible among all the available clusters (See Algorithm 2). Every time a bad decision is made, some important elements might be blocked; namely, they won’t be selected in future iterations. So, there will be some covered and some blocked elements at every iteration. Figure 2 exemplifies this situation, where only $OPT/2$ elements are covered by the returned feasible solution. In fact, $1/2$ is the approximation ratio of this algorithm (See Theorem 1, which relies on Lemmas 1, 2, and 3); Figure 2 serves as a tight example.
@@
1:$p$ clusters $\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{p}$
2:A set $S$ containing one subset from each cluster
3:$S,C\leftarrow\emptyset$
4:$K\leftarrow\{1,2,...,p\}$
5:while $K\not=\emptyset$ do
6:     $k\leftarrow\operatorname*{argmax}_{j\in K}\max_{\mathcal{S}\in\mathcal{C}_{j}}|\mathcal{S}-C|$
7:     $K\leftarrow K-\{k\}$
8:     $A\leftarrow\operatorname*{argmax}_{\mathcal{S}\in\mathcal{C}_{k}}|\mathcal{S}-C|$
@@
Figure 2: In this instance for CMCP, $p=6$, $\mathcal{C}_{1}=\{\mathcal{S}_{1,1},\mathcal{S}_{1,2}\}$, $\mathcal{C}_{2}=\{\mathcal{S}_{2,1},\mathcal{S}_{2,2}\}$, $\mathcal{C}_{3}=\{\mathcal{S}_{3,1},\mathcal{S}_{3,2}\}$, $\mathcal{C}_{4}=\{\mathcal{S}_{4,1}\}$, $\mathcal{C}_{5}=\{\mathcal{S}_{5,1}\}$, and $\mathcal{C}_{6}=\{\mathcal{S}_{6,1}\}$. The optimal solution $\{\mathcal{S}_{1,1},\mathcal{S}_{2,1},\mathcal{S}_{3,1},\mathcal{S}_{4,1},\mathcal{S}_{5,1},\mathcal{S}_{6,1}\}$ covers $OPT=30$ squares. If the greedy approximation algorithm selects $\mathcal{S}_{1,2}$ at the first iteration, all the squares inside $\mathcal{S}_{1,1}$ get blocked because no other subsets from the remaining clusters cover them. If the algorithm selects $\mathcal{S}_{2,2}$ in the second iteration, all the squares inside $\mathcal{S}_{2,1}$ get blocked. Finally, if the algorithm selects $\mathcal{S}_{3,2}$ at the third iteration, all the squares inside $\mathcal{S}_{3,1}$ get blocked. Thus, the next three selected subsets won’t be able to cover any uncovered squares and the returned solution will be $\{\mathcal{S}_{1,2},\mathcal{S}_{2,2},\mathcal{S}_{3,2},\mathcal{S}_{4,1},\mathcal{S}_{5,1},\mathcal{S}_{6,1}\}$, which covers $OPT/2=15$ squares.
@@
At each iteration of Algorithm 2, the number of newly blocked elements cannot exceed the number of newly covered elements. By “newly” we mean that those elements were not blocked or covered in previous iterations.
@@
Let $\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{p}$ be the clusters of the input instance. Let $S^{*}=\{\mathcal{S}_{1}^{*},\mathcal{S}_{2}^{*},...,\mathcal{S}_{p}^{*}\}$ be an optimal solution and $S^{\prime}=\{\mathcal{S}_{1}^{\prime},\mathcal{S}_{2}^{\prime},...,\mathcal{S}_{p}^{\prime}\}$ be the feasible solution returned by Algorithm 2, where $\mathcal{S}_{i}^{*},\mathcal{S}_{i}^{\prime}\in\mathcal{C}_{i}$. A set $\mathcal{S}_{j}^{\prime}$ with maximum coverage $|\mathcal{S}_{j}^{\prime}-C|$ is selected at every iteration, where $C$ is the set of covered elements at previous iterations. If $\mathcal{S}_{j}^{\prime}\not\in S^{*}$ (or equivalently, if $\mathcal{S}_{j}^{\prime}\not=\mathcal{S}_{j}^{*}$), then in the worst scenario all the elements in $\mathcal{S}_{j}^{*}-C$ are not in the subsets from the remaining clusters; therefore, they get blocked. Since $|\mathcal{S}_{j}^{\prime}-C|$ is maximal, $|\mathcal{S}_{j}^{*}-C|\leq|\mathcal{S}_{j}^{\prime}-C|$. ∎
Lemma 2 shows how the size of the optimal solution for a given instance $\Pi_{\alpha}$ is related to the size of the optimal solution of smaller instances $\Pi_{\beta}$ that result from removing some cluster and some specific elements from $\Pi_{\alpha}$. This lemma will be useful to prove the next propositions.
Lemma 2.
Let $\Pi_{\alpha}$ be an arbitrary instance for the GBP with clusters $\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{p}$, an optimal solution size of $OPT(\Pi_{\alpha})$, and an optimal solution $S^{*}=\{\mathcal{S}_{1}^{*},\mathcal{S}_{2}^{*},...,\mathcal{S}_{p}^{*}\}$, where $\mathcal{S}_{1}^{*}\in\mathcal{C}_{1}$, $\mathcal{S}_{2}^{*}\in\mathcal{C}_{2}$, and so on. Let $\mathcal{S}_{a}$ be an arbitrary subset from an arbitrary cluster $\mathcal{C}_{a}$ and let $\mathcal{S}_{a}^{*}$ be the one element in $\mathcal{C}_{a}\cap S^{*}$. Besides, let $\Pi_{\beta}$ be an instance constructed as follows: it has the same clusters as $\Pi_{\alpha}$, except $\mathcal{C}_{a}$, and $\mathcal{S}_{a}\cap\cup_{\{\mathcal{S}\in\mathcal{C}_{j}:j\in[1,p]-\{a\}\wedge\mathcal{S}\in S^{*}\}}\mathcal{S}$ is removed from the universe of elements. The optimal solution size of $\Pi_{\beta}$ is $OPT(\Pi_{\alpha})-y-c$, where $y$ is the number of elements only $\mathcal{S}_{a}^{*}$ can cover (these would be the blocked elements in Algorithm 2), i.e.,
$y=|\mathcal{S}_{a}^{*}-\mathcal{S}_{a}^{*}\cap\cup_{\{\mathcal{S}\in\mathcal{C}_{j}:j\in[1,p]-\{a\}\wedge\mathcal{S}\in S^{*}\}}\mathcal{S}|\ \ ,$ (18)
and $c$ is the number of elements in $\mathcal{S}_{a}$ that would contribute to $OPT(\Pi_{\beta})$ if they were not removed, i.e.,
$c=|\mathcal{S}_{a}\cap\cup_{\{\mathcal{S}\in\mathcal{C}_{j}:j\in[1,p]-\{a\}\wedge\mathcal{S}\in S^{*}\}}\mathcal{S}|\ \ .$ (19)
Proof.
Let $\Pi_{\beta}^{\prime}$ be the instance that results by only removing $\mathcal{C}_{a}$ from $\Pi_{\alpha}$. Notice that $OPT(\Pi_{\beta}^{\prime})$ cannot be smaller than $OPT(\Pi_{\alpha})-y$ because this is the number of elements covered by all the subsets in $S^{*}-\{\mathcal{S}_{a}^{*}\}$. Thus, $OPT(\Pi_{\beta}^{\prime})\geq OPT(\Pi_{\alpha})-y$. In case $OPT(\Pi_{\beta}^{\prime})>OPT(\Pi_{\alpha})-y$, then $OPT(\Pi_{\alpha})<OPT(\Pi_{\beta}^{\prime})+y$, which contradicts the optimality of $OPT(\Pi_{\alpha})$. Therefore, $OPT(\Pi_{\beta}^{\prime})=OPT(\Pi_{\alpha})-y$. By a similar argument, $OPT(\Pi_{\beta})$ cannot be different from $OPT(\Pi_{\beta}^{\prime})-c$. Therefore, $OPT(\Pi_{\beta})=OPT(\Pi_{\beta}^{\prime})-c=OPT(\Pi_{\alpha})-y-c$. ∎
To better understand Lemma 2, consider an instance $\Pi_{\alpha}$ with $p=4$, $\mathcal{C}_{1}=\{\mathcal{S}_{1,1},\mathcal{S}_{1,2}\}$, $\mathcal{C}_{2}=\{\mathcal{S}_{2,1},\mathcal{S}_{2,2}\}$, $\mathcal{C}_{3}=\{\mathcal{S}_{3,1}\}$, and $\mathcal{C}_{4}=\{\mathcal{S}_{4,1}\}$; it has an optimal solution $\{\mathcal{S}_{1,1},\mathcal{S}_{2,1},\mathcal{S}_{3,1},\mathcal{S}_{4,1}\}$ and $OPT(\Pi_{\alpha})=10$ (See Figure 3).
Figure 3: An instance $\Pi_{\alpha}$ with $p=4$ and $OPT(\Pi_{\alpha})=10$.
By selecting $\mathcal{S}_{2,2}$ as the “arbitrary” subset from the “arbitrary” cluster $\mathcal{C}_{2}$, removing $\mathcal{C}_{2}$ from $\Pi_{\alpha}$ results in $\Pi_{\beta}^{\prime}$, which has $p=3$, $\mathcal{C}_{1}=\{\mathcal{S}_{1,1},\mathcal{S}_{1,2}\}$, $\mathcal{C}_{3}=\{\mathcal{S}_{3,1}\}$, and $\mathcal{C}_{4}=\{\mathcal{S}_{4,1}\}$ (See Figure 4); it has an optimal solution $\{\mathcal{S}_{1,1},\mathcal{S}_{3,1},\mathcal{S}_{4,1}\}$ and $OPT(\Pi_{\beta}^{\prime})=OPT(\Pi_{\alpha})-y=10-1=9$, where $y=1$ is the number of blocked elements, i.e., elements in $\mathcal{S}_{2,1}-\mathcal{S}_{2,1}\cap(\mathcal{S}_{1,1}\cup\mathcal{S}_{3,1}\cup\mathcal{S}_{4,1})$ (See Eq. (18)).
Figure 4: An instance $\Pi_{\beta}^{\prime}$ with $p=3$ and $OPT(\Pi_{\beta}^{\prime})=9$. $\Pi_{\beta}^{\prime}$ results from removing $\mathcal{C}_{2}$ from $\Pi_{\alpha}$.
Afterward, by removing the elements in $S_{2,2}$ that contribute to $OPT(\Pi_{\beta}^{\prime})$ from $\Pi_{\beta}^{\prime}$, we get $\Pi_{\beta}$ (See Figure 5). Thus, $OPT(\Pi_{\beta})=OPT(\Pi_{\beta}^{\prime})-c=OPT(\Pi_{\alpha})-y-c=10-1-2=7$, where $c$ is the cardinality of $\mathcal{S}_{2,2}\cap(\mathcal{S}_{1,1}\cup\mathcal{S}_{3,1}\cup\mathcal{S}_{4,1})$ (See Eq. (19)).
Figure 5: An instance $\Pi_{\beta}$ with $p=3$ and $OPT(\Pi_{\beta})=7$. $\Pi_{\beta}$ results from removing some specific elements from $\Pi_{\beta}^{\prime}$.
Algorithm 2 iteratively prunes the input instance by removing a specific cluster; hence the relevance of Lemma 2. In more detail, by the end of every iteration $i$ of Algorithm 2, an instance $\Pi_{i}$ with optimal solution size $OPT(\Pi_{i})$ is obtained, being $\Pi_{0}$ the original input instance. Each $\Pi_{i}$ results from pruning $\Pi_{i-1}$ as follows: remove the cluster $\mathcal{C}_{l_{i}}$ that contains the largest subset $\mathcal{S}_{l_{i}}$ among all available clusters (line 5) and remove all elements in $\mathcal{S}_{l_{i}}$ from the universe of elements (lines 4, 6, and 8 of Algorithm 2). Using this notation, a lower bound on $OPT(\Pi_{i})$ is derived in the next lemma.
Lemma 3.
$z_{i}=z_{i-1}-y_{i}-x_{i}$ is a lower bound for $OPT(\Pi_{i})$, where $y_{i}$ is the number of newly blocked elements at iteration $i$, $x_{i}$ is the number of newly covered elements at iteration $i$, and $z_{0}=OPT(\Pi_{0})$.
Proof.
By definition,
$z_{i}=z_{i-1}-y_{i}-x_{i}\ \ .$ (20)
By Lemma 2,
$OPT(\Pi_{i})=OPT(\Pi_{i-1})-y_{i}-c_{i}\ \ ,$ (21)
where $y_{i}$ is the number of newly blocked elements (See Eq. (18)) and $c_{i}$ is the cardinality of a subset of $\mathcal{S}_{l_{i}}-C$ (See Eq. (19)). Therefore,
$c_{i}\leq|\mathcal{S}_{l_{i}}-C|=x_{i}\ \ ,$ (22)
where $C$ is the set of covered elements at previous iterations. Notice that keeping track of $C$ at lines 4, 6, and 8 of Algorithm 2 is equivalent to removing all the previously covered elements from each $\Pi_{i}$. Since $z_{0}=OPT(\Pi_{0})$, Eq. (21) leads to
$OPT(\Pi_{1})=z_{0}-y_{1}-c_{1}\ \ .$ (23)
By Eq. (22), $c_{1}\leq x_{1}$, which along with Eq. (20) and Eq. (23) leads to
$z_{1}=z_{0}-y_{1}-x_{1}\leq z_{0}-y_{1}-c_{1}=OPT(\Pi_{1})\ \ .$ (24)
Therefore,
$z_{1}\leq OPT(\Pi_{1})\ \ .$ (25)
For $i=2$, Eq. (20) and Eq. (25) lead to
$z_{2}=z_{1}-y_{2}-x_{2}\leq OPT(\Pi_{1})-y_{2}-x_{2}\ \ .$ (26)
Then, Eq. (26), Eq. (21) and $c_{2}-x_{2}\leq 0$ (Eq. (22)) lead to
$z_{2}\leq OPT(\Pi_{2})+c_{2}-x_{2}\leq OPT(\Pi_{2})\ \ .$ (27)
Thus,
$z_{2}\leq OPT(\Pi_{2})\ \ .$ (28)
For $i\in[3,p]$, the pattern continues:
$z_{i}=z_{i-1}-y_{i}-x_{i}\leq OPT(\Pi_{i-1})-y_{i}-x_{i}\ \ ,$ (29)
$z_{i}\leq OPT(\Pi_{i})+c_{i}-x_{i}\leq OPT(\Pi_{i})\ \ ,$ (30)
$z_{i}\leq OPT(\Pi_{i})\ \ .$ (31)
@@
The greedy approximation algorithm for CMCP (Algorithm 2) has an approximation ratio of $1/2$.
@@
Let $\Pi_{0}$ be an input instance for the GBP. Let $OPT(\Pi_{0})$ be the size of the optimal solution, let $x_{i}$ be the number of newly covered elements at iteration $i$, and let $y_{i}$ be the number of newly blocked elements at iteration $i$. Since $x_{1}$ is maximal,
$x_{1}\geq\frac{OPT(\Pi_{0})}{p}\ \ \ \ ;$ (32)
otherwise, $px_{1}<OPT(\Pi_{0})$, which means that $p$ disjoint largest subsets would not be able to cover $OPT(\Pi_{0})$ elements, which is a contradiction. Using Lemma 3, this argument can be generalized to all iterations $i\in[1,p]$ in terms of $z_{i}$,
$x_{i}\geq\frac{OPT(\Pi_{i-1})}{p-(i-1)}\geq\frac{z_{i-1}}{p-(i-1)}\ \ \ \ ,$ (33)
Notice that $z_{0}=OPT(\Pi_{0})$ and $z_{p}=0$ because all $z_{p-1}$ vertices are covered at the last iteration. Besides, by Lemma 1, no more than $x_{i}$ elements can be blocked at each iteration:
$y_{i}\leq x_{i}\ \ \ \ .$ (34)
With these observations, the theorem is proved by contradiction. Let us assume that the algorithm returns a solution that covers less than $OPT(\Pi_{0})/2$ elements, i.e.,
$\sum_{i=1}^{p}x_{i}<\frac{OPT(\Pi_{0})}{2}\ \ \ \ .$ (35)
Eq. (34) and Eq. (35) lead to
$\sum_{i=1}^{p}y_{i}\leq\sum_{i=1}^{p}x_{i}<\frac{OPT(\Pi_{0})}{2}\ \ \ \ .$ (36)
Then, Eq. (20) and Eq. (36) lead to
$\sum_{i=1}^{p}z_{i}>\sum_{i=1}^{p}z_{i-1}-OPT(\Pi_{0})\ \ \ \ .$ (37)
Notice that $\sum_{i=1}^{p}z_{i-1}-\sum_{i=1}^{p}z_{i}=z_{0}-z_{p}$. Since $z_{0}=OPT(\Pi_{0})$ and $z_{p}=0$, Eq. (37) reduces to
$OPT(\Pi_{0})=z_{0}-z_{p}<OPT(\Pi_{0})\ \ \ \ ,$ (38)
which is a contradiction. Therefore, $\sum_{i=1}^{p}x_{i}\geq OPT(\Pi_{0})/2$. Namely, the greedy approximation algorithm for CMCP has an approximation ratio of $1/2$.
@@ 4.2 A greedy heuristic for GBP
Algorithm 3 shows the proposed deterministic greedy heuristic (Gr) for the GBP, which implements the greedy approximation algorithm for CMCP. In a nutshell, starting from the clusters with larger subsets, i.e., from $p-1$ to 0 (line 4), Gr iteratively selects the corresponding vertex of the subset that covers as many uncovered elements as possible (lines 8 and 9); notice that ‘$+$’ adds a vertex at the right side of the sequence, which is initially empty. In more detail, Gr receives as input a graph $G=(V,E)$, a guess $p$ on $b(G)$, and the distance matrix $D$, which can be computed independently using breadth-first search. Then, Gr computes the closed neighborhood $N_{p-1}[v]$ of each $v\in V$. In other words, it computes the adjacency list of $G^{p-1}$ (line 3). Afterward, it iteratively selects $p$ vertices that correspond to the subsets that cover as many uncovered vertices as possible in graphs $G^{r}$, where $r$ goes from $p-1$ to 00 (lines 4 to 11). Notice that the adjacency list is updated at each iteration, containing uncovered vertices at a distance at most $r$ from each $v\in V$ (line 6). Besides, every subset $N_{i}[v]$ of each cluster is a superset of each $N_{j}[v]$, where $j<i$. Thus, the corresponding vertex of a subset with maximum coverage among all the clusters is selected at each iteration. Therefore, this is a correct implementation of Algorithm 2. Since the distance matrix $D$ is known, updating the adjacency list can be done in at most $\mathcal{O}(n^{2})$ steps at each iteration. Thus, the overall complexity of Gr is $\mathcal{O}(pn^{2})$; however, if we consider the time required to compute all-pairs’ shortest path, its complexity is $\mathcal{O}(mn+pn^{2})$ as stated in the abstract. By Theorem 1, Proposition 3 follows.
@@
1:A graph $G=(V,E)$, its distance matrix $D$, and a guess $p$ on $b(G)$
2:A sequence $S$
3:$S\leftarrow()$
4:$B\leftarrow\emptyset$
5:$N_{p-1}\leftarrow$ Adjacency list of $G^{p-1}$
6:for $r=p-1$ to 00 do
7:     if $r<p-1$ then
8:         $N_{r}\leftarrow$ Update the adjacency list for $G^{r}$
9:     end if
10:     $v\leftarrow\operatorname*{argmax}_{u\in V}|N_{r}[u]-B|$
11:     $S\leftarrow S+v$
12:     $B\leftarrow B\cup N_{r}[v]$
13:end for
14:return $S$
@@
If the subsets Gr selects cover less than $n/2$ vertices, then $p<b(G)$.
@@
By Theorem 1, Algorithm 2 returns a solution that covers at least $OPT/2$ elements, where $OPT$ is the optimal solution size. By definition, a burning sequence burns $n$ vertices. So, if $p\geq b(G)$, $OPT=n$, and since Gr correctly implements Algorithm 2, it must return a solution that covers at least $n/2$ vertices. By contrapositive, if Gr returns a solution that covers less than $n/2$ vertices, then $p<b(G)$. ∎
For Gr to be useful, we must know $b(G)$. We can try different guesses $p$ to tackle the issue of not knowing this value in advance. The simpler way is by trying all $p\in[1,n]$. However, a binary search is more efficient. Although Proposition 3 helps find a lower bound for $b(G)$, we didn’t find it useful in practice. Therefore we didn’t integrate it into the binary search (See Algorithm 4), which we now explain. First, the distance between every pair of vertices is computed in $\mathcal{O}(mn+n^{2})$ steps using breadth-first search (line 1). Then, the lower and upper bounds are computed using the burning farthest-first (BFF) approximation algorithm for GBP [10], which runs in $\mathcal{O}(n^{2})$ steps if the computation of all-pairs’ shortest path is not considered; Algorithm 5 shows a detailed pseudocode for BFF using a first-in-first-out (FIFO) queue. Since BFF returns a solution $S_{0}$ (line 2) of length at most $23b(G)-$ 2 [10], the lower bound on $b(G)$ can be safely set to $\lceil(s(S_{0})+2)/3\rceil$ (line 3) and the upper bound to $s(S_{0})-1$ (line 4), where $s(S_{0})$ is the size of $S_{0}$. Then, Gr is executed with parameters $G$, $p$, and $D$ (lines 6 and 7); GrP will be introduced later. Afterward, every time Gr returns a burning sequence, $h$ is set to $p-1$; otherwise, $l$ is set to $p+1$. This way, Gr is executed $\mathcal{O}(\log n)$ times with different guesses on $p$ that might approach $b(G)$. Using Gr at line 7, and since $p\leq n$, the complexity of Algorithm 4 is $\mathcal{O}(n^{3}\log n)$.
@@
5:$l\leftarrow\lceil(s(S_{0})+2)/3\rceil$
6:$h\leftarrow s(S_{0})-1$
@@
8:     $p\leftarrow\lfloor(l+h)/2\rfloor$
9:     $S\leftarrow$ Gr$(G,p,D)$ or GrP$(G,p,D)$
@@
2:A burning sequence $S$ of length at most $23b(G)-$ 2
@@
5:$S\leftarrow(v)$
6:$q.push(v)$
7:$B\leftarrow\{v\}$
8:for $u\in V$ do
9:     $dist(u)\leftarrow d(u,v)$
10:end for
11:while $B\not=V$ do
12:     $q\_size\leftarrow q.size()$
13:     for $j=1$ to $q\_size$ do
14:         $v\leftarrow q.pop()$
15:         for $u\in N(v)-B$ do
16:              $q.push(u)$
17:              $B\leftarrow B\cup\{u\}$
18:         end for
@@
20:     $v\leftarrow\operatorname*{argmax}_{u\in V}dist(u)$
21:     $S\leftarrow S+v$
22:     $q.push(v)$
23:     $B\leftarrow B\cup\{v\}$
24:     for $u\in V$ do
25:         if $d(u,v)<dist(u)$ then
26:              $dist(u)\leftarrow d(u,v)$
27:         end if
28:     end for
29:end while
30:return $S$